4. Angelo was choosing springs to use their door. He noticed that spring A was stiffer than spring B; that is ka> ka. Which spring will have more elastic potential energy stored if (a) they are compressed the same distance and (b) if the same force is applied on them?







1
Expert's answer
2021-11-11T12:00:47-0500

Given,

Spring constant of sprint A "=K_a"

Spring constant of spring B "=K_b"

As we know that spring A is stiffer than spring B, so "K_a>K_b"

If same force(F) applied on both the spring then the elongation in the spring be "x_a" and "x_b"

"x_a=\frac{F}{K_a}" and "x_b=\frac{F}{K_b}"

Now, potential energy stored in the spring

"E_a=\frac{k_a F^2}{2K_a^2}"


"E_b=\frac{k_b F^2}{2K_b^2}"


Now, "\frac{E_a}{E_b}=\frac{k_b}{k_a}"